# A Comprehensive Review on Harnessing Large Language Models to Overcome Recommender System Challenges

### v1

@@ 5. Limitations & Open Challenges of LLMs in Recommender Systems
Latency and Throughput Bottlenecks – Autoregressive decoding imposes sequential delays that violate the sub–100 ms SLA common in web-scale ranking. Even with 4-bit weight-only quantization, a 7-B model exceeds 35 ms per query on a single A100 (Rajaraman et al., 2025). Token-parallel kernels, KV-cache fusion, and Flash-Attention v3 yield 1.7×1.7\!$\times$1.7 ×2.3×2.3\!$\times$2.3 × speed-ups but degrade under bursts that saturate memory bandwidth (Jo et al., 2025; Shah et al., 2024). Hierarchical routing—first-stage dense retrieval, second-stage sparse MoE—cuts tail latency by 28 % yet adds routing overhead and fairness concerns (Tairin et al., 2025). On-device NPUs and speculative decoding shrink $\approx\!40$ ms, but require tight integration with vector search pre-rankers to avoid head-of-line blocking (Lee et al., 2024; Yang et al., 2025).
@@
Operational Cost and Resource Footprint – Inference FLOPs scale roughly linearly with prompt length and quadratically with hidden width; doubling model size can quadruple serving cost and 3× carbon footprint (Zhang et al., 2023). Sparse MoE routing and LoRA adapters reduce average compute by 60–80 % but necessitate elastic load balancing and suffer from prompt skew (Li et al., 2025; Jin et al., 2025). Token-level caching amortizes 15–25 % of compute but introduces cache consistency bugs when prompts embed personalization tokens (Guo et al., 2025). Workload-aware autoscaling and serverless accelerators lower idle GPU burn but add cold-start penalties that hurt tail latency (Tran et al., 2024). Cloud API mark-ups (\$$\sim\$0.03$ per k-tokens) further inflate total cost of ownership in high-throughput pipelines (Rogers et al., 2025).
@@
Hallucination and Semantic Misalignment – LLMs can fabricate item attributes, user intents, or reasoning chains, surfacing non-existent products or biased analogies (Surana and Thompson, 2023; Patel and Deshmukh, 2024). Retrieval grounding, constrained decoding, and post-generation fact-checkers cut hallucinations by 25–40 % but remain brittle on long-tail, code-mixed, or multilingual content  (Yu et al., 2024; Chen et al., 2025b; Gao et al., 2024). Fidelity critics trained with human feedback reduce error rate another 12 ms latency per call (Shao et al., 2024). Negative prompt augmentation and contrastive rejection sampling further reduce hallucination probability to $<4\%$ on internal ad-catalog tests (Liu and Zhu, 2025). Nonetheless, hallucinations that slip through can propagate via feedback loops, degrading long-term engagement metrics (Pan et al., 2024).
@@
Prompt Sensitivity and Reproducibility – Microscopic lexical edits (e.g., “recommend” → “suggest”) can produce ranking deltas $>\!10$ NDCG, complicating A/B parity, rollback, and incident triage (Savarese et al., 2020; Mitchell et al., 2019). Prompt provenance graphs, canary tokens, and contrastive diff testing improve reproducibility but incur  5 GB/day in lineage logs per million requests (Zhuo et al., 2024). RL-based prompt search raises robustness by 18 % but sacrifices transparency and locks systems into brittle reward hacks (Park and James, 2025). Hash-based prompt fingerprints enable fast cache look-ups but fail under adversarial paraphrasing (Liu and Zhao, 2025). A standardized robustness benchmark for recsys prompts is still missing (Zhu et al., 2024).
@@
Representation Drift and Embedding Incompatibility – Periodic fine-tunes or RLHF updates shift latent spaces, invalidating historical user vectors and cached scores (McNee et al., 2006). Embedding versioning, backward-compatible projection heads, and dual-encoder rehearsal maintain cosine similarity $>\!0.85$ across upgrades, but add 2–3 ms per inference and triple index storage (Liu and Rehn, 2025; Qian et al., 2024). Continual-learning regularizers slow drift but sacrifice 1–2 % relevance on fresh items (Zhao and Meier, 2025). Online distillation into lightweight twin-towers shows promise yet fails for zero-shot entities (Kim and Vento, 2025). Drift alarms based on Wasserstein distance flag latent shifts, but trigger frequent false positives under seasonal traffic variation (Xu and Khan, 2024).
@@
Evaluation Gaps – Classic list metrics (NDCG, MAP, Recall@K) ignore semantic fidelity, novelty, and explanation quality central to LLM outputs. Hybrid metrics that blend counterfactual click uplift with LLM-judged coherence scores correlate weakly with real satisfaction ($r\!<\!0.35$(Joachims et al., 2018; Yin and Huang, 2025). LLM-as-judge pipelines reduce annotation cost 10× but inherit biases from the same model family (Bougie and Watanabe, 2025; Wataoka et al., 2024). Causal bandit simulators yield better offline-online fidelity but require calibrated reward scaling and realism audits (Nie and Li, 2024). Large-scale human-in-the-loop testing remains cost-prohibitive, averaging \$0.42 per labeled explanation (Tschiatschek et al., 2024).

### v2

@@ 5. Limitations & Open Challenges of LLMs in Recommender Systems
Latency and Throughput Bottlenecks – Autoregressive decoding imposes sequential delays that violate the sub–100 ms SLA common in web-scale ranking. Even with 4-bit weight-only quantization, a 7-B model exceeds 35 ms per query on a single A100 (Rajaraman et al., 2025). Token-parallel kernels, KV-cache fusion, and Flash-Attention v3 yield $1.7\!\times$$2.3\!\times$ speed-ups but degrade under bursts that saturate memory bandwidth (Jo et al., 2025; Shah et al., 2024). Hierarchical routing—first-stage dense retrieval, second-stage sparse MoE—cuts tail latency by 28 % yet adds routing overhead and fairness concerns (Tairin et al., 2025). On-device NPUs and speculative decoding shrink $\approx\!40$ ms, but require tight integration with vector search pre-rankers to avoid head-of-line blocking (Lee et al., 2024; Yang et al., 2025).
@@
Operational Cost and Resource Footprint – Inference FLOPs scale roughly linearly with prompt length and quadratically with hidden width; doubling model size can quadruple serving cost and 3× carbon footprint (Zhang et al., 2023). Sparse MoE routing and LoRA adapters reduce average compute by 60–80 % but necessitate elastic load balancing and suffer from prompt skew (Li et al., 2025; Jin et al., 2025). Token-level caching amortizes 15–25 % of compute but introduces cache consistency bugs when prompts embed personalization tokens (Guo et al., 2025). Workload-aware autoscaling and serverless accelerators lower idle GPU burn but add cold-start penalties that hurt tail latency (Tran et al., 2024). Cloud API mark-ups ($\sim\mathdollar 0.03$ per k-tokens) further inflate total cost of ownership in high-throughput pipelines (Rogers et al., 2025).
@@
Hallucination and Semantic Misalignment – LLMs can fabricate item attributes, user intents, or reasoning chains, surfacing non-existent products or biased analogies (Surana and Thompson, 2023; Patel and Deshmukh, 2024). Retrieval grounding, constrained decoding, and post-generation fact-checkers cut hallucinations by 25–40 % but remain brittle on long-tail, code-mixed, or multilingual content  (Yu et al., 2024; Chen et al., 2025b; Gao et al., 2024). Fidelity critics trained with human feedback reduce error rate another 12 ms latency per call (Shao et al., 2024). Negative prompt augmentation and contrastive rejection sampling further reduce hallucination probability to $4\%$ on internal ad-catalog tests (Liu and Zhu, 2025). Nonetheless, hallucinations that slip through can propagate via feedback loops, degrading long-term engagement metrics (Pan et al., 2024).
@@
Prompt Sensitivity and Reproducibility – Microscopic lexical edits (e.g., “recommend” → “suggest”) can produce ranking deltas >10>\!10 % NDCG, complicating A/B parity, rollback, and incident triage (Savarese et al., 2020; Mitchell et al., 2019). Prompt provenance graphs, canary tokens, and contrastive diff testing improve reproducibility but incur  5 GB/day in lineage logs per million requests (Zhuo et al., 2024). RL-based prompt search raises robustness by 18 % but sacrifices transparency and locks systems into brittle reward hacks (Park and James, 2025). Hash-based prompt fingerprints enable fast cache look-ups but fail under adversarial paraphrasing (Liu and Zhao, 2025). A standardized robustness benchmark for recsys prompts is still missing (Zhu et al., 2024).
@@
Representation Drift and Embedding Incompatibility – Periodic fine-tunes or RLHF updates shift latent spaces, invalidating historical user vectors and cached scores (McNee et al., 2006). Embedding versioning, backward-compatible projection heads, and dual-encoder rehearsal maintain cosine similarity >0.85>\!0.85 across upgrades, but add 2–3 ms per inference and triple index storage (Liu and Rehn, 2025; Qian et al., 2024). Continual-learning regularizers slow drift but sacrifice 1–2 % relevance on fresh items (Zhao and Meier, 2025). Online distillation into lightweight twin-towers shows promise yet fails for zero-shot entities (Kim and Vento, 2025). Drift alarms based on Wasserstein distance flag latent shifts, but trigger frequent false positives under seasonal traffic variation (Xu and Khan, 2024).
@@
Evaluation Gaps – Classic list metrics (NDCG, MAP, Recall@K) ignore semantic fidelity, novelty, and explanation quality central to LLM outputs. Hybrid metrics that blend counterfactual click uplift with LLM-judged coherence scores correlate weakly with real satisfaction (r<0.35r\!<\!0.35(Joachims et al., 2018; Yin and Huang, 2025). LLM-as-judge pipelines reduce annotation cost 10× but inherit biases from the same model family (Bougie and Watanabe, 2025; Wataoka et al., 2024). Causal bandit simulators yield better offline-online fidelity but require calibrated reward scaling and realism audits (Nie and Li, 2024). Large-scale human-in-the-loop testing remains cost-prohibitive, averaging \$0.42 per labeled explanation (Tschiatschek et al., 2024).